# Conformal e-prediction

### v3

@@ 10 Conclusion
In this paper we have discussed (even if some of them very briefly) four strengths of conformal prediction:
@@
    In the online prediction protocol, conformal predictors as set predictors make errors at different steps independently. (Conformal e-predictors satisfy a weakened version of this property, as discussed in Sect. 3.)
@@
    In the online prediction protocol, conformal prediction produces bona fide test martingales, which can then be used in one-off and multistage conformal testing. (Weakened variants for conformal e-prediction are discussed in Sections 8 and 9.)
  4. 4.
@@
Two of these strengths, 1 and 4, appear to be clear advantages of conformal prediction over conformal e-prediction. For strengths 2 and 3, this is less obvious, but the picture for conformal prediction still appears simpler and nicer.
@@
  • Designing conditional conformal e-predictors is much easier and using e-values adds flexibility, especially as compared with Mondrian conformal predictors; see Sect. 4.
  • Cross-conformal e-predictors are provably valid, unlike cross-conformal predictors, as discussed in Sect. 6.
Looking for further advantages is an interesting direction of further research.
@@ Acknowledgments
Many thanks to Aaditya Ramdas for pointing out serious deficiencies in presentation in an earlier version. This research has been partially supported by Astra Zeneca, Stena Line, and Mitie.
@@ Appendix A Law of the iterated logarithm for e-flows
Let $\mathcal{F}_{n}$, where $n\in\mathbb{Z}$ and $\mathbb{Z}$ stands for the set of integer numbers, be an increasing sequence of $\sigma$-algebras on a given sample space $\Omega$,
@@
where $\mathcal{F}_{n}$, $n=1,2,\dots$, is the exchangeable filtration (introduced in Sect. 3: $\mathcal{F}_{n}$ is generated by $\lbag Z_{1},\dots,Z_{n-1}\rbag,Z_{n},Z_{n+1},\dots$), and
@@
where $E_{n}$ is the $n$th conformal e-variable, $E_{n}:=f(Z_{1},\dots,Z_{n})$. Let me call this the conformal e-flow. The fact that the adapted sequence $(\xi_{n},\mathcal{G}_{n})$ defined in this way is an e-flow is the main property of validity of conformal e-prediction. The main part of the conformal e-flow is $(\xi_{n},\mathcal{G}_{n})$ for $n<0$; the extension to all integer $n$ is vacuous. Exact e-flows correspond to admissible conformal predictors.
There are two natural laws of the iterated logarithm for e-flows: the forward law describes the behaviour, as $N\to\infty$, of the sums $\sum_{n=1}^{N}(\xi_{n}-1)$, and the backward law describes the behaviour of the sums $\sum_{n=-N}^{-1}(\xi_{n}-1)$ The forward law is just a restatement of the standard martingale law of the iterated logarithm, since with each forward sequence $\xi_{n}$, $n\geq 1$, we can associate a supermartingale that carries the same information by setting
$X_{n}:=\sum_{i=1}^{n}(\xi_{i}-1),$ (32)
with $X_{0}:=0$; remember that the defining property of being a supermartingale is $\mathbb{E}(X_{n}\mid\mathcal{F}_{n-1})=X_{n-1}$. Applying the standard law of the iterated logarithm to the supermartingale $X_{n}$ gives us the following corollary.
@@
Such a reduction to the supermartingale case is impossible for the backward law. Luckily, however, the standard proof of the law of the iterated logarithm can be easily adapted to the backward case and allows us to establish the following backward law of the iterated logarithm.
@@
The proof of Proposition 22 uses the standard basic scheme (see, e.g., [21, Sect. 4.4]). As a first step we fix an e-flow $(\xi_{n},\mathcal{G}_{n})$ and assume, without loss of generality, that it is exact (an easy way to see that there is no loss of generality is to apply the idea of coupling [4, Sect. 7.4]).
We need an auxiliary exponential supermartingale from Stout’s proof of the law of the iterated logarithm for martingales [22, Lemma 5.4.1].
@@
Therefore, we can choose
$\psi(n)\sim\sqrt{2c\lambda n\ln\ln n}.$
The conditions $\delta\leq 1/c$ and $\delta\leq 1$ mentioned earlier indeed hold, from some $k$ on, for (37), $\gamma:=k^{\lambda}$, and $N:=n_{k}$.
@@
Appendix B Proofs for Sect. 9
The idea of the proof of Proposition 19 consists, as usual, in reversing the direction of time. Let $N$ be a sufficiently large natural number, and let $\mathcal{F}_{n}$, $n\in\{1,\dots,N+1\}$, be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and observations $Z_{n},\dots,Z_{N}$. We also allow $n=N+1$, in which case $\mathcal{F}_{N+1}$ is the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{N}\rbag$. This is a finite-horizon version of the exchangeable filtration.
Let $E_{n}$ be the $n$th conformal e-variable (having $e_{n}$ as it value). Then $(E_{n},\mathcal{F}_{n})$, $n=N,\dots,1$, is an exact e-flow over the finite time interval $N,\dots,1$, in the sense
$\mathbb{E}(E_{n}\mid\mathcal{F}_{n+1})=1,\quad n=N,\dots,1.$
The corresponding martingale is $(T_{n},\mathcal{F}_{n})$, $n=N+1,\dots,1$, where
$T_{n}:=E_{n}\dots E_{N},\quad n=N+1,N,\dots,1,$
with $T_{N+1}$ understood to be $1$. (But now we compound $E_{n}$ in a multiplicative manner, while in the previous appendix, e.g., in (32), compounding was additive.) For simplicity, let us assume that all $E_{n}$ are positive, so that $T_{n}$ is a positive martingale.
Let us apply the Shiryaev–Roberts procedure to the e-flow $E_{N},\dots,E_{1},\dots$ continued by setting $E_{0}:=E_{-1}:=\dots:=1$. It gives us the decreasing sequence of stopping times $\sigma_{0}:=N+1$ and
$\sigma_{k}:=\max\left\{n<\sigma_{k-1}\mid\sum_{i=n}^{\sigma_{k-1}-1}E_{n}\dots E_{i}\geq c\right\},\quad k=1,2,\dots,$ (42)
where $n$ ranges over the integers. A useful property of the Shiryaev–Roberts procedure is $\mathbb{E}(\sigma_{k-1}-\sigma_{k}\mid\mathcal{F}_{\sigma_{k}})\geq c$; see, e.g., [28, Proposition 8.13] (proved in [28, Sect. 8.5.2], while the current proof uses other ideas in that section).
To make the stopping times (42) more manageable, it will be convenient to force an alarm every $L$ steps, where $L$ is to be chosen later: $\sigma^{\prime}_{0}:=N+1$ and
$\sigma^{\prime}_{k}:=(\sigma^{\prime}_{k-1}-L)\vee\max\left\{n<\tau_{k-1}\mid\sum_{i=n}^{\sigma^{\prime}_{k-1}-1}E_{n}\dots E_{i}\geq c\right\},\quad k=1,2,\dots;$
let us call this the truncated Shiryaev–Roberts procedure. As in the proof of [28, Corollary 8.4], induction in $k$ shows that $\sigma^{\prime}_{k}\geq\sigma_{k}$ for all $k$.
There is a useful connection between $\tau_{k}$ and $\sigma^{\prime}_{k}$: namely, each set $\{\tau_{k-1}+1,\dots,\tau_{k}\}$ with $\tau_{k}\leq N$ contains at least one stopping time $\sigma^{\prime}_{l}$. (This is even true for $\sigma_{l}$, although we do not need it.) This can be deduced from
$\sum_{j=i}^{\tau_{k}}E_{i}\dots E_{j}\geq c\text{ for some $i\geq\tau_{k-1}+1$}.$ (43)
Indeed, let $l$ be the largest number satisfying $\sigma^{\prime}_{l}>\tau_{k}$; arguing indirectly, let us suppose that $\sigma^{\prime}_{l+1}<\tau_{k-1}+1$. The inequality (43) implies
$\sum_{j=i}^{\sigma^{\prime}_{l}-1}E_{i}\dots E_{j}\geq\sum_{j=i}^{\tau_{k}}E_{i}\dots E_{j}\geq c\text{ for some $i>\sigma^{\prime}_{l+1}$},$
which contradicts the definition of the stopping times $\sigma^{\prime}$.
As in [28, Proof of Proposition 8.15 in Sect. 8.5.2], let us say that $k$ is slow if
$\mathbb{P}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}=L\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c/L$
and fast otherwise. We show there that
$\mathbb{E}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c-c^{2}/L$ (44)
if $k$ is fast. On the other hand, if $k$ is slow,
$\mathbb{E}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq L\mathbb{P}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}=L\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c.$
In both cases, we have (44).
Let $A^{\prime}_{N}$ be the largest $k$ such that $\sigma^{\prime}_{k}>0$; we interpret $A^{\prime}_{N}$ as the number of alarms raised by the truncated Shiryaev–Roberts procedure. Remember that $A_{N}$, defined by (28), is the number of alarms raised by the conformal reverse Shiryaev–Roberts e/procedure. By Hoeffding’s inequality [28, Sect. A.6.3], for arbitrarily small $\epsilon\in(0,c)$ and sufficiently large $N$,
$\mathbb{P}\left(\frac{A_{N}}{N}\geq\frac{1}{c-\epsilon}\right)$ $\leq\mathbb{P}\left(\frac{A^{\prime}_{N}}{N}\geq\frac{1}{c-\epsilon}\right)=\mathbb{P}\left(A^{\prime}_{N}\geq\left\lceil\frac{N}{c-\epsilon}\right\rceil\right)$
$\leq\mathbb{P}\left(\sum_{k=1}^{\lceil\frac{N}{c-\epsilon}\rceil}(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k})\leq N\right)$
$\leq\exp\left(-2\frac{\left(\lceil\frac{N}{c-\epsilon}\rceil\left(c-\frac{c^{2}}{L}\right)-N\right)^{2}}{L^{2}\lceil\frac{N}{c-\epsilon}\rceil}\right)\leq\exp(-\epsilon^{\prime}N),$
where $\epsilon^{\prime}$ is a positive constant (which requires a sufficiently large $L$). Since the series $\sum_{N}\exp(-\epsilon^{\prime}N)$ converges,
$\frac{A_{N}}{N}\geq\frac{1}{c-\epsilon}$
happens only finitely often, which completes the proof of Proposition 19.
To deduce Proposition 18 notice that, by induction in $k$, $\tau_{k}^{\text{SR}}\leq\tau_{k}^{\text{CUSUM}}$, where $\tau_{k}^{\text{CUSUM}}$ are the $\tau_{k}$ defined by (27), and $\tau_{k}^{\text{SR}}$ are the $\tau_{k}$ defined by (30).

### v4

@@ 7 Conclusion
In this paper we have discussed three strengths of conformal prediction (even if two of them briefly):
@@
    In the online prediction protocol, conformal predictors as set predictors make errors at different steps independently. Conformal e-predictors satisfy a weakened version of this property, as discussed in Sect. 2.
@@
Two of these strengths, 1 and 3, appear to be clear advantages of conformal prediction over conformal e-prediction. For strength 2, this is less obvious, but the picture for conformal prediction still appears simpler and nicer.
@@
  • Designing conditional conformal e-predictors is much easier and using e-values adds flexibility, especially as compared with Mondrian conformal predictors; see Sect. 3.
  • Cross-conformal e-predictors are provably valid, unlike cross-conformal predictors, as discussed in Sect. 5.
Looking for further advantages of conformal e-prediction is an interesting direction of further research.
@@ Acknowledgments
Many thanks to Aaditya Ramdas for pointing out deficiencies in presentation in an earlier version and to Alex Balinsky and Ilia Nouretdinov for useful discussions. This research has been partially supported by Astra Zeneca, Stena Line, and Mitie.
@@ Appendix A Law of the iterated logarithm for e-flows
Let $\mathcal{F}_{n}$, where $n\in\mathbb{Z}$ and $\mathbb{Z}$ stands for the set of all integer numbers, be an increasing sequence of $\sigma$-algebras on a given sample space $\Omega$,
@@
where $\mathcal{F}_{n}$, $n=1,2,\dots$, is the exchangeable filtration (introduced in Sect. 2: $\mathcal{F}_{n}$ is generated by $\lbag Z_{1},\dots,Z_{n-1}\rbag,Z_{n},Z_{n+1},\dots$), and
@@
where $E_{n}$ is the $n$th conformal e-variable, $E_{n}:=f(Z_{1},\dots,Z_{n})$. Let us call this the conformal e-flow. The fact that the adapted sequence $(\xi_{n},\mathcal{G}_{n})$ defined in this way is an e-flow is the main property of validity of conformal e-prediction. The main part of the conformal e-flow is $(\xi_{n},\mathcal{G}_{n})$ for $n<0$; the extension to all integer $n$ is vacuous. Exact e-flows correspond to admissible conformal predictors.
There are two natural laws of the iterated logarithm for e-flows: the forward law describes the behaviour, as $N\to\infty$, of the sums $\sum_{n=1}^{N}(\xi_{n}-1)$, and the backward law describes the behaviour of the sums $\sum_{n=-N}^{-1}(\xi_{n}-1)$ The forward law is just a restatement of the standard martingale law of the iterated logarithm, since with each forward sequence $\xi_{n}$, $n\geq 1$, we can associate a supermartingale $X$ that carries the same information by setting
$X_{n}:=\sum_{i=1}^{n}(\xi_{i}-1),$
with $X_{0}:=0$; remember that the defining property of being a supermartingale is $\mathbb{E}(X_{n}\mid\mathcal{F}_{n-1})\leq X_{n-1}$. Applying the standard law of the iterated logarithm to the supermartingale $X_{n}$ gives us the following corollary.
@@
Such a reduction to the supermartingale case is impossible for the backward law; even though there are numerous laws of the iterated logarithm for reverse martingales, they are not applicable in our current context. Luckily, however, the standard proof of the law of the iterated logarithm can be easily adapted to the backward case and allows us to establish the following backward law of the iterated logarithm.
@@
The proof of Proposition 17 uses the standard basic scheme (see, e.g., [19, Sect. 4.4]). As a first step we fix an e-flow $(\xi_{n},\mathcal{G}_{n})$ and assume, without loss of generality, that it is exact (an easy way to see that there is no loss of generality is to apply the idea of coupling [4, Sect. 7.4]).
We need an auxiliary exponential supermartingale from Stout’s proof of the law of the iterated logarithm for martingales [20, Lemma 5.4.1].
@@
Therefore, we can choose $\psi(n)\sim\sqrt{2c\lambda n\ln\ln n}$. The conditions $\delta\leq 1/c$ and $\delta\leq 1$ mentioned earlier indeed hold, from some $k$ on, for (32), $\gamma:=k^{\lambda}$, and $N:=n_{k}$.
@@
Appendix B Bounding the error probability of conformal e-predictors via Markov’s inequality
The first advantage of conformal prediction mentioned in Sect. Conformal e-prediction is that conformal predictors can be used as set predictors, in which case their property of validity can be expressed as a low probability of error. In principle, this can also be done in the case of conformal e-prediction, and has been done in [1, Sect. 2], but it leads to a predictor that is not admissible in the terminology of statistical decision theory [29, Sect. 1.3]. This is the topic of this appendix. For simplicity we will assume that the object and label spaces $\mathbf{X}$ and $\mathbf{Y}$ are finite.
Let us fix the size $n$ of the training set. A set predictor is a function $\Gamma:\mathbf{Z}^{n}\times\mathbf{X}\to 2^{\mathbf{Y}}$. It is $\epsilon$-valid, where $\epsilon>0$, if $\mathbb{P}(Y\notin\Gamma(Z_{1},\dots,Z_{n},X))\leq\epsilon$ provided $Z_{1},\dots,Z_{n},(X,Y)$ are exchangeable; in other words, if the probability of error is bounded by $\epsilon$. An example of an $\epsilon$-valid set predictor is the conformal $\epsilon$-predictor
$\Gamma^{\epsilon}(z_{1},\dots,z_{n},x):=\{y:f(z_{1},\dots,z_{n},x,y)>\epsilon\},$
where $f$ is a conformal p-predictor [25, Proposition 2.3]. Conformal $\epsilon$-predictors are just conformal predictors packaged as set predictors.
The BB-predictor [1, Sect. 2] associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is defined as
$\Gamma(z_{1},\dots,z_{n},x):=\left\{y:f(z_{1},\dots,z_{n},x,y)<1/\epsilon\right\}.$
In combination with Markov’s inequality, Proposition 1 implies that the BB-predictor defined in this way is $\epsilon$-valid.
An $\epsilon$-valid set predictor $\Gamma$ is inadmissible if there exists an $\epsilon$-valid set predictor $\Gamma^{\prime}$ such that
$\Gamma^{\prime}(z_{1},\dots,z_{n},x)\subseteq\Gamma(z_{1},\dots,z_{n},x)$ (37)
for all $z_{1},\dots,z_{n},x$ and the inclusion is strict for some $z_{1},\dots,z_{n},x$. Otherwise, $\Gamma$ is admissible. We will say that $\Gamma^{\prime}$ dominates $\Gamma$ if (37) holds for all $z_{1},\dots,z_{n},x$ and that the domination is strict if the inclusion in (37) is strict for some $z_{1},\dots,z_{n},x$.
First let us check that each BB-predictor is dominated by a conformal predictor.
Proposition 21.
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is $\epsilon$-valid, but it is dominated by the conformal $\epsilon$-predictor based on $f$’s nonconformity e-measure.
Proof.
The validity was checked earlier. Let us check that the BB-predictor is dominated by the conformal $\epsilon$-predictor $\Gamma^{\epsilon}$ (based on $f$’s nonconformity e-measure as conformity measure): if
|$\frac{\left|\{i=1,\dots,n+1:\alpha_{i}\geq\alpha_{n+1}\}\right|}{n+1}>\epsilon,$
then we have $\alpha_{i}\geq\alpha_{n+1}$ for at least $\lfloor(n+1)\epsilon\rfloor+1$ $\alpha_{i}$s, which implies
$\frac{\alpha_{n+1}}{\frac{1}{n+1}\sum_{i=1}^{n+1}\alpha_{i}}\leq\frac{n+1}{\lfloor(n+1)\epsilon\rfloor+1}<\frac{1}{\epsilon}.\qed$
In typical cases a BB-predictor will be inadmissible, being strictly dominated by the conformal predictor based on the same nonconformity (e-)measure. This will formalized in the next proposition, for which we need two definitions.
A nonconformity e-measure $A$ is generic if it always outputs $(\alpha_{1},\dots,\alpha_{n+1})$ that are all different (they may be only slightly different, or alternatively we can add slight randomization). In this case we will also say that the corresponding conformal e-predictor is generic.
A nonconformity measure $A$ is $\epsilon$-categorical if its nonconformity scores take only two values, 00 and $1/\epsilon$: it only outputs $(\alpha_{1},\dots,\alpha_{n+1})$ with $\alpha_{i}\in\{0,1/\epsilon\}$ for all $i\in\{1,\dots,n+1\}$. Categorical nonconformity e-measures $A$ are a way of encoding set predictors: such an $A$ represents the set predictor
$\Gamma(z_{1},\dots,z_{n},x):=\left\{y\in\mathbf{Y}:f(z_{1},\dots,z_{n},x,y)>0\right\},$
where $f$ is the corresponding conformal e-predictor. The next proposition says that the only way to avoid inadmissibility of the BB-predictor based on a generic nonconformity e-measure $A$ is to make $A$ very close to being $\epsilon$-categorical, where the deviation from being $\epsilon$-categorical is measured by
$d(A):=\max_{(z_{1},\dots,z_{n+1})\in\mathbf{Z}^{n+1}}\left(\sum_{i:\alpha_{i}\geq 1/\epsilon}(\alpha_{i}-1/\epsilon)+\sum_{i:\alpha_{i}<1/\epsilon}\alpha_{i}\right),$
$\alpha_{1},\dots,\alpha_{n+1}$ are the nonconformity scores for $z_{1},\dots,z_{n+1}$, and $i$ ranges over $\{1,\dots,n+1\}$. We will also say that $d(f):=d(A)$ is the deviation from being $\epsilon$-categorical for the conformal e-predictor $f$ associated with $A$.
Proposition 22.
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is inadmissible if $f$ is generic and $d(f)\geq 1/\epsilon$.
The lower bound of $1/\epsilon$ on the deviation $d(f)$ in Proposition 22 is very small for a large size $n$ of the training set, as the typical order of magnitude for $d(f)$ is $n$. Therefore, BB-predictors are typically inadmissible.
Proof of Proposition 22.
In order for the BB-predictor to be admissible, for each $(z_{1},\dots,z_{n+1})\in\mathbf{Z}^{n+1}$, the $\lfloor(n+1)\epsilon\rfloor$ largest $\alpha_{i}$ in the corresponding $(\alpha_{1},\dots,\alpha_{n+1})$ should all be at least $1/\epsilon$. Therefore, the deviation from being $\epsilon$-categorical should be at most
$n+1-\frac{\lfloor(n+1)\epsilon\rfloor}{\epsilon}<n+1-\frac{(n+1)\epsilon-1}{\epsilon}=1/\epsilon.\qed$